1. Use metric conversion factors to solve each of the following problems a. A Bee Hummingbird has a mass of 0.0018 kg. What is the mass of the hummingbird in micrograms (ug)? b. An assay protocol requires you to add 12.5 mg of lysozyme to 1.0 L of buffer. You have a lysozyme stock vial labeled 25 mg/mL. How many microliters (μL) of the stock will you need to add to your solution. Express 0.5 mg/mL in μg/uL Express 0.4 to mmol/mL C. d.
Answer #1

a.

1000 g 1 kg 10° , 1 1.8 × 106 μg 0.0018 kg × × .

b.

\frac{12.5 \ mg}{25 \ mg/mL}=0.50 \ mL

0.50 \ mL \times \frac{10^3 \ \mu L}{1 \ mL}=5.0 \times 10^2 \ \mu L

c.

\frac{0.5 \ mg}{1 \ mL} \times \frac {1 \ mL}{1000 \ \mu L} \times \frac {1000 \ \mu L} {1 \ mL}=0.5 \ \mu g/ \mu L

d.

0.4 \ \mu M = \frac{0.4 \ \mu mol}{1 \ L}

= \frac{0.4 \ \mu mol}{1 \ L} \times \frac{1 \ L}{1000 \ mL} \times \frac{1 \ mmol}{1000 \ \mu mol} =4 \times 10^{-7} \ mmol/mL